OpenStudy (anonymous):

PLEASE HELP!!! Write a equation in slope-intercept form for an equation of a line perpendicular to y=2x+4 and that intersects at point (4,3)

OpenStudy (helder_edwin):

what is the slope of the line u r given?

OpenStudy (anonymous):

I have no clue sorry

OpenStudy (helder_edwin):

your equation has the form \[ \large y=mx+b \] here m is the slope and b is the y-intercept

OpenStudy (helder_edwin):

this means m=2 and the line passes through (0,4) right?

OpenStudy (anonymous):

i have the 4 choices of a anwser but thats all tht came with that ?

OpenStudy (anonymous):

yea

OpenStudy (helder_edwin):

now when you have two lines say \[ \large y=m_1x+b \] \[ \large y=m_2x+c \] for them to be perpendicular it must be \[ \large m_1\cdot m_2=-1 \] you have m1=2 find m2

OpenStudy (anonymous):

kay i got the rest thanks